The following table gives the grouped data on the weights of all 100 babies born at Clinic last year.

Weight (Units) Number of babies

3 to less than 5 5

5 to less than 7 30

7 to less than 9 40

9 to less than 11 20

11 to less than 13 5

(a) Calculate the mean

(b) Calculate the variance and standard deviation

(c) Calculate the mode

(d) Calculate the median

(e) Calculate the coefficient of variation


1
Expert's answer
2021-09-15T02:57:52-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 Class & f & \ \ x\ \ & \ f(x)\ & f( x^2) & cf\\ 
\hline
 3-5 & 5 & 4 & 20 & 80 & 5\\
 \hdashline
 5-7 & 30 & 6 & 180 & 1080 & 35\\
 \hdashline
 7-9 & 40 & 8 & 320 & 2560 & 75\\
 \hdashline
 9-11 & 20 & 10 & 200 & 2000 & 95\\
 \hdashline
 11-13 & 5 & 12 & 60 & 720 & 100\\
 \hdashline
 & n & & \sum f (x) & \sum f 


(x^2) & & \\
 & =100 & & =780 & =6440 & \\
\end{array}"

(a)


"mean=\bar{x}=\dfrac{\sum_if_i(x_i)}{n}=\dfrac{780}{100}=7.8"

(b)


"\sigma^2=\dfrac{\sum_if_i(x_i)^2-(\sum_if_i(x_i))^2}{n}"

"=\dfrac{6440-(780)^2}{100}=3.56"

"\sigma=\sqrt{3.56}\approx1.8868"

(c)

Maximum frequency is 40. The mode class is 7-9.

"L =" lower boundary point of mode class "=7"

"f_1="frequency of the mode class "=40"

"f_0="  frequency of the preceding class "=30"

"f_2=" frequency of the succedding class "=20"

"c="  class length of mode class "=2"


"mode=Z=L+(\dfrac{f_1-f_0}{2f_1-f_0-f_2})\cdot c"

"=7+(\dfrac{40-30}{2(40)-30-20})\cdot2\approx7.6667"

(d)

value of "(n\/2)th"  observation "=" value of "(100\/2)th"  observation "="

value of "(50)th"  observation

From the column of cumulative frequency "cf," we find that the "(50)th"  observation lies in the class 7-9.

The median class is 7-9.

"L=" lower boundary point of median class "=7"

"n=" Total frequency "=100"

"cf=" Cumulative frequency of the class preceding the median class "=35"

"f=" Frequency of the median class "=40"

"c=" class length of median class "=2"


"median=M=L+(\dfrac{\dfrac{n}{2}-cf}{f})\cdot c"

"=7+(\dfrac{50-35}{40})\cdot 2=7.75"

(e)

coefficient of variation "=\dfrac{\sigma}{\bar{x}}\cdot100\%=\dfrac{1.8868}{7.8}\cdot100\%\approx24.19\%"